find the area under the standard normal curve to the left of z= -2.52 and to the right of z= -0.58
Answer #1

(a) z < -2.52

The probability that Z<−2.52 is equal to the blue area under the curve.

P ( Z<−2.52 ) = 1−P ( Z < 2.52 )

P ( Z<−2.52 ) = 1−0.9941

P ( Z<−2.52) =0.0059

(b) Z > -0.58

P ( Z>−0.58 )=P ( Z<0.58 )

P ( Z>−0.58 ) = P ( Z<0.58 ) = 0.719


P ( Z>−0.58 ) = 0.719